### v1

@@ 2.2.2 Escape criterion
When planets are in resonances and under the stable criterion, the resonant quantities ($\phi$,$\Phi$,$\eta$) remain fixed with a small amplitude liberation. Here we attempt to answer what occurs if the above stability criterion is not satisfied. We define such a state as overstablity. In this case, the perturbations grow exponentially over time (no longer in the linear regime). Whether the overstability leads planets directly escape or still stay in resonance can be further analyzed by the phase space topology.
@@
Figure 4 shows the Hamiltonian curves for different $\eta$ values in the phase space of the conjugate variables $X{=}\sqrt{\Phi}\cos{\varphi}$ and $Y{=}\sqrt{\Phi}\sin{\varphi}$. We note that $\eta$ represents the deviation from the exact resonance and the distance to the origin corresponds to the excited eccentricity strength ($\sqrt{X^{2}+Y^{2}}{=}\sqrt{\Phi}{\sim}e$).
@@
Importantly, the value of $\eta$ sets the number of fixed points, and therefore determines the subsequent dynamical evolution. The fixed points can be obtained from $\partial H/\partial\varphi{=}\partial H/\partial\Phi{=}0$ (also see Eq.8-135 of Murray & Dermott 1999). It can be shown that, for $\eta{>}-3/2$, there is only one stable fixed point (black dot in Figure 4), whereas for $\eta{<}-3/2$, there are two stable fixed points and one unstable point (red dot in Figure 4).
@@
In the overstable circumstances, the perturbation near the equilibrium grows significantly. When the equilibrium value $\eta_{\rm eq}{>}-3/2$, the system’s Hamiltonian curve expands as eccentricities excited during resonant capture. However, the trajectory still bounds to the only fixed point. As a result, the system’s eccentricity, period ratio, and resonant angle librate near the equilibrium. Following Goldreich & Schlichting (2014) and Xu & Lai (2017), we refer to this state as an overstable trap, which is also equivalent to circumstance described as the limit cycle in Afkanpour et al. (2024). One example of overstable trap can be found in Figure 3.
@@
The separatrix occurs when $\eta_{\rm eq}{<}-3/2$ and the topology completely changes, as shown in Figure 4. In this case, the system’s Hamiltonian curve originally sits near the right stable fixed point. As the curve expands, it cross the separatrix and enter the circulation zone centered around the left stable fixed point (from the right to the left black dot in Figure 4). After then, the curve continuously shrinks towards the left black fixed point. As such, the eccentricity decreases, and the weaker resonant interaction cannot keep the planet pair in resonance anymore. The planets’ period ratio departs from the exact commensurability, and the system eventually escapes from the resonance. One example of escape is shown in Figure 5.
Based on Eq. (19c), $\eta_{\rm eq}{<}-3/2$ requires $\Phi_{\rm eq}{=}(j-1)\beta\tau_{\rm e}/(4j\tau_{a}){>}1$, which is corresponding to
$\frac{\tau_{a}}{\tau_{e}}<\left(\frac{3}{f_{\rm d,i}}\right)^{2/3}\frac{(j-1)^{8/9}j^{4/9}}{8j}\mu_{\rm o}^{-2/3}\sim\mu_{\rm o}^{-2/3}.$ (23)
@@
It is worth noting that this escape condition requires the system is already in the overstable regime. In the case where the mass of the outer planet is much lower than that of the inner one ($q\gg 1$), the resonance is always stable. As such, the escape does not occur even when $\eta_{\rm eq}{<}-3/2$.
Taking together, resonance trapping is always stable when the mass of the inner planet is much higher than that of the outer one ($q{\gg}1$). However, when the mass of the outer planet is much higher than that of the inner one ($q{\ll}1$), the outcome can be classified into three configurations: stable trap, overstable trap, or escape. We further extend the analysis to arbitrary mass ratios $q$ and derive the corresponding criteria as follows:
$\begin{cases}{\displaystyle\left(\frac{\tau_{a}}{\tau_{e}}\right)_{\rm overstab}=\left(\frac{3}{\mu_{o}}\right)^{2/3}h(q)\ \ \left[\frac{(j-1)^{2}}{f_{\rm d,i}\alpha}\right]^{2/3}(j-1)^{2/3}f(q)}\\ {\displaystyle}\hfill[\mbox{stability criterion}],\vspace{0.1cm}\\ {\displaystyle\left(\frac{\tau_{a}}{\tau_{e}}\right)_{\rm esc}=\left(\frac{3}{\mu_{o}}\right)^{2/3}\frac{h(q)}{4}\ \ \left[\frac{(j-1)^{2}+j^{2}q}{f_{\rm d,i}\alpha+f_{\rm d,o}^{{}^{\prime}}q^{2}}\right]^{2/3}}\\ {\displaystyle}\hfill[\mbox{escape criterion}].\end{cases}$ (24)
@@
$h(q)$ $=$ $\frac{1}{2}\frac{1}{1+q\sqrt{\alpha}}\frac{f_{\rm d,i}^{2}+f_{\rm d,o}^{{}^{\prime}2}q^{2}\alpha}{jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}},$ (25)
@@
Mention that $h(q)=\tau_{a}/\tau_{e}(e_{\rm i,eq}^{2}+e_{\rm o,eq}^{2})$, $f(q)=8(j-1)[1-(e_{\rm o,eq}/e_{\rm i,eq})^{2}]$. Eq. (24) reduces to Eqs. (22) and (23) in the limit $q\ll 1$. We show the detailed derivation in the Appendix B.
Figure 5: Resonance crossing of a two-planet system with convergent migration. The initial starting from the outer-to-inner planet period ratio is $2.2$ (slightly beyond the $2$:$1$ MMR), and migration parameters are $\tau_{a}=2.5\times 10^{5}$ and $\tau_{a}/\tau_{e}=10^{2}$. The period ratio between the inner and outer planets exhibits small oscillations upon entering resonance and ultimately decreases further away from the 2:1 ratio. The eccentricity of the inner planet undergoes significant oscillations around the resonant equilibrium value after resonance is established, eventually damping to a very small value. The resonant angle of the inner planet maintains large oscillations upon entering resonance and eventually diverges. In phase space, the inner planet initially oscillates around the resonant fixed point, but the oscillation range increases until it is ultimately captured by the origin.
Figure 6: The theoretical distribution of the four kinds of evolution near 2:1 MMR depending on different $\tau_{a}$ and $\tau_{a}/\tau_{e}$. When $\tau_{a}$ and $\tau_{a}/\tau_{e}$ are in the blue region, the two planets pass through the resonance because the eccentricity damping is too strong. When $\tau_{a}$ and $\tau_{a}/\tau_{e}$ are in the blown region, they can’t be trapped into the resonance since they migrate inward too fast. The green region indicates that the planets can maintain a stable resonance. The yellow area represents that although the resonance trapping is overstable, the planets can still maintain a 2:1 period ratio. The purple region indicates that the planets will eventually leave the 2:1 MMR position and migrate inward after overstability. The left panel shows the distribution when the outer planet is more massive than the inner one, and the right panel displays the distribution when the inner planet is more massive than the outer one.
@@ 2.3 A unified picture
We summarize different outcomes considering the dynamics evolution of planet pairs that undergo convergent migration. The illustrated runs from the numerical simulations in each regime are also presented.
  1. No trap (Figure 1): The system cannot reach the resonant equilibrium state and the planet pair directly bypasses the resonance. The planets’ eccentricities increase and their period ratio temporarily holds during the resonant passage. However, after rapid crossing the resonance, the eccentricity drops and the resonant angle varies from 00 to $2\pi$.
@@
  2. Stable trap (Figure 2): The planets are locked in resonance with statically stable orbits. The planets’ period ratio, eccentricities, and resonant angle remain at the equilibrium values after capture. Meanwhile, the trajectory of the system converges towards a fix point in the phase space.
@@
  3. Overstable trap (Figure 3): The planets are locked in resonance with dynamically stable orbits. Their eccentricities and resonance angle initially converge to the equilibrium values. After a while, the eccentricities and resonance angle liberate with increasing amplitudes. This state is nevertheless stable and the period ratio of planets remains constant after capture. In the phase space, the system’s trajectory initially approaches the fixed point, then begins to oscillate, and ultimately settles into a limit cycle around that point.
@@
  4. Escape (Figure 5): The planets are locked in resonance with only temporarily stable orbits. Their period ratio, eccentricities and resonance angle initially converge to the equilibrium values. The planets’ eccentricities and resonance angle quickly oscillate with increasing amplitudes. This state becomes unstable and the planets eventually leave resonance. As such, the planets’ eccentricities decrease, and the period ratio continues to drop, and the resonant angle changes from liberation to circulation. In the phase space, the system’s trajectory initially approaches the fixed point; however, the oscillation gradually grows until it intersects the separatrix. The trajectory ultimately settles towards the coordinate origin.
We point out that although several literature studies have explored this issue in depth (Huang & Ormel, 2023; Goldreich & Schlichting, 2014; Deck & Batygin, 2015; Batygin, 2015), none have formulated a comprehensive model. Here we demonstrate a unified picture in Figure 6. The outcome is given in the $\tau_{a}$-$\tau_{e}$ parameter space. We describe how $\tau_{\rm a}$ and $\tau_{\rm e}$ influence the outcomes as follows.
@@
If convergent migration is too rapid, the planets directly bypass the resonance without reaching the equilibrium state. As migration slows down, the effect of eccentricity damping becomes significant. Strong e-damping has two consequences: the planet resonant interaction gets weaken and semimajor axis decays due to orbital circularization. When the orbital decay surpasses the resonant term, the planets will no longer remain in resonance. In this regard, planets cannot get capture when $\tau_{\rm e}$ and/or $\tau_{\rm a}$ are short. No trap occurs in the upper and left part of Figure 6.
@@
When planets are trapped in resonance, the stability is related to the degree of eccentricity excitation. The damping effect ($\tau_{\rm e}$) reduces the planet’s eccentricity, leading the system stable. On the other hand, convergent migration ($\tau_{\rm a}$) brings planets closer together, in turn exciting their eccentricities. Consequently, the relative eccentricity damping $\tau_{\rm a}/\tau_{\rm e}$ is the key parameter affecting the stability. As can be seen from top to bottom of Figure 6, system tends to be more unstable as $\tau_{\rm a}/\tau_{\rm e}$ decreases.

### v2

@@ 2.2.2 Escape criterion
When planets are in resonances and under the stable criterion the resonant quantities ($\phi$$\Phi$$\eta$) remain fixed with a small amplitude libration. Here we attempt to answer what occurs if the above stability criterion is not satisfied. This state is commonly described as overstability (Goldreich & Schlichting, 2014; Deck & Batygin, 2015; Xu & Lai, 2017). In this case, the perturbations grow exponentially over time due to non-adiabatic effects (eccentricity damping in this case). Whether the overstability leads planets to escape directly or still stay in resonance can be further analyzed by the phase space topology.
@@
Figure 4 shows the Hamiltonian curves for different $\eta$ values in the phase space of the conjugate variables $X{=}\sqrt{\Phi}\cos{\varphi}$ and $Y{=}\sqrt{\Phi}\sin{\varphi}$. We note that $\eta$ represents the deviation from the exact resonance and the distance to the origin corresponds to the excited eccentricity ($\sqrt{X^{2}+Y^{2}}{=}\sqrt{\Phi}{\sim}e$).
@@
Importantly, the value of $\eta$ sets the number of fixed points, and therefore determines the subsequent dynamical evolution. The fixed points can be obtained from $\partial H/\partial\varphi{=}\partial H/\partial\Phi{=}0$. It can be shown that (Ferraz-Mello, 2007; Batygin, 2015; Petit et al., 2017), for $\eta{>}-3/2$, there is only one stable fixed point (black dot in Figure 4), whereas for $\eta{<}-3/2$, there are two stable fixed points and one unstable point (red dot in Figure 4).
@@
In overstable circumstances, the perturbations near the equilibrium grow over time. When the equilibrium value $\eta_{\rm eq}{>}-3/2$, the system’s Hamiltonian curve expands as eccentricities are excited during resonant capture. However, the trajectory still bounds to the only fixed point. As a result, the system’s eccentricity, period ratio, and resonant angle librate around their equilibrium values. Following Goldreich & Schlichting (2014) and Xu & Lai (2017), we refer to this state as an overstable trap – although the equilibrium state is overstable and perturbations grow over time ($\delta\Phi\propto e^{\lambda_{\rm r}t}$, $\lambda_{\rm r}{>}0$), the system remains trapped with the resonant angle librating at a finite amplitude. This state is also termed as the limit cycle in Deck & Batygin (2015); Nesvorný et al. (2022); Afkanpour et al. (2024). One example of an overstable trap can be found in Figure 3.
@@
The separatrix emerges when $\eta_{\rm eq}{<}-3/2$ and the topology completely changes, as shown in Figure 4. In this case, the system’s Hamiltonian curve originally sits near the right stable fixed point. As the curve expands, it crosses the separatrix and enters the circulation zone centered around the left stable fixed point (from the right to the left black dot in Figure 4). After that, the curve continuously shrinks towards the left black fixed point. As such, the eccentricity decreases, and the weaker resonant interaction cannot keep the planet pair in resonance anymore. The planets’ period ratio departs from the exact commensurability, and the system eventually escapes from the resonance. One example of escape is shown in Figure 5.
Based on Eq. (19c), $\eta_{\rm eq}{<}-3/2$ requires $\Phi_{\rm eq}{=}(j-1)\beta\tau_{e}/(2j\tau_{\rm m}){>}1$, which is corresponding to
$\frac{\tau_{\rm m}}{\tau_{e}}<\left(\frac{3}{\mu_{\rm o}}\right)^{2/3}\frac{1}{4j}\left[\frac{(j-1)^{2}}{f_{\rm d,i}\alpha}\right]^{2/3}\sim\mu_{\rm o}^{-2/3}.$ (23)
@@
It is worth noting that this escape condition requires the system already to be in the overstable regime. In the case where the mass of the outer planet is much lower than that of the inner one ($q\gg 1$), the resonance is always stable. As such, the escape does not occur even when $\eta_{\rm eq}{<}-3/2$.
To summarise, resonance trapping is always stable when the mass of the inner planet is much higher than that of the outer one ($q{\gg}1$). However, when the mass of the outer planet is much higher than that of the inner one ($q{\ll}1$), the outcome can be classified into three configurations: stable trap, overstable trap, or escape. We further extend the analysis to arbitrary mass ratios $q$ and derive the corresponding criteria as follows:
$\begin{cases}{\displaystyle\frac{\tau_{\rm m}}{\tau_{e}}>\left(\frac{3}{\mu_{o}}\right)^{2/3}h(q)\ \ \left[\frac{(j-1)}{f_{\rm d,i}\alpha}\right]^{2/3}f(q)}\\ {\displaystyle}\hfill[\mbox{stability criterion}],\vskip 2.84544pt\\ {\displaystyle\frac{\tau_{\rm m}}{\tau_{e}}<\left(\frac{3}{\mu_{o}}\right)^{2/3}\frac{h(q)}{4}\ \ \left[\frac{(j-1)^{2}+j^{2}q}{f_{\rm d,i}\alpha+f_{\rm d,o}^{{}^{\prime}}q^{2}}\right]^{2/3}}\\ {\displaystyle}\hfill[\mbox{escape criterion}],\end{cases}$ (24)
@@
$h(q)$ $=$ $\frac{1}{(1+q\sqrt{\alpha})}\frac{f_{\rm d,i}^{2}+f_{\rm d,o}^{{}^{\prime}2}q^{2}\alpha}{jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}},$ (25)
@@
We note that $h(q)$ and $f(q)$ can also be written as $h(q){=}\tau_{\rm m}/\tau_{e}(e_{\rm i,eq}^{2}+e_{\rm o,eq}^{2})$, and $f(q){=}1-(e_{\rm o,eq}/e_{\rm i,eq})^{2}$. In the limit $q{\ll}1$, Eq. (24) is equivalent to Eqs. (22) and (23) . We show the detailed derivation in Appendix B.
Figure 5: Escape from resonance in a two-planet system with convergent migration. The planet and disk parameters are $m_{\rm i}{=}1\ M_{\oplus}$, $m_{\rm o}{=}10\ M_{\oplus}$, $\tau_{\rm m}{=}5\times 10^{5}$ yr and $\tau_{\rm m}/\tau_{e}{=}200$. The system is temporarily trapped in a 22:11 MMR during $t{\sim}2{-3}\times 10^{4}$ yr, and the inner planet reaches the equilibrium eccentricity with growing amplitude. The system escapes from resonance afterward, with $e_{\rm i}$ quickly damping to a very low value. Similarly, $\varphi_{\rm i}$ features an increasing libration amplitude during the trap and eventually oscillates from 0 to $2\pi$. In phase space, the trajectory initially librates around the right fixed point (black dot). As the amplitude grows, it ultimately escapes from the fixed point and settles into the origin. See https://github.com/llh-astro/resonance/raw/main/gif/escape.gif
Figure 6: Analytical predictions of two-planet systems undergoing convergent migration near the 22:11 MMR in the limits of $q{\ll}1$ (left) and $q{\gg}1$ (right). Blue region: planets directly bypass the MMR due to strong eccentricity damping. Brown region: convergent migration is too fast to allow resonance capture. Green region: planets get trapped in a stable resonance. Yellow region: despite overstability occurs, planets still get trapped in the resonance, with libration of finite amplitude around the equilibrium point. Purple region: after a temporary trap, planets eventually escape resonance and continue the inward migration. The illustrative examples of the above regimes can be found in Figure 151{-}5.
@@ 2.3 A unified picture
We summarize the resonance capture outcomes for planet pairs that undergo convergent migration. Representative simulation results from each regime are also presented.
  1. No trap (Figure 2): The planets directly bypasses the resonance. After crossing the resonance, the planets’ eccentricities drop and the resonant angles vary from 0 to $2\pi$.
@@
  2. Stable trap (Figure 2): The planets get locked in resonance, and the system’s trajectory converges towards an equilibrium value in the phase space. The planets’ period ratio, eccentricities, and resonant angles remain at equilibrium values after capture.
@@
  3. Overstable trap (Figure 3): The planets are locked in resonance with librating resonant angles. The eccentricities and the resonance angles of planets initially converge to the equilibrium values and, after a while, librate with increasing amplitudes. The period ratio of the planets remains constant after capture. In the phase space, the system’s trajectory initially approaches the fixed point, then transitions into oscillations before ultimately stabilizing into a limit cycle around that point.
@@
  4. Escape (Figure 5): The planets are temporarily locked in resonance before eventually escaping. The period ratio, eccentricities, and resonance angle of the planet pair initially converge to the equilibrium values. Then the eccentricities and resonance angles start to oscillate with growing amplitudes. Eventually, the planets leave resonance, with a shift in resonant angles from libration to circulation. In phase space, the system’s trajectory initially approaches the fixed point; however, the oscillation gradually amplifies until it crosses the separatrix. The trajectory ultimately stabilizes towards the coordinate origin.
We point out that several previous studies have explored this issue in depth (Huang & Ormel, 2023; Goldreich & Schlichting, 2014; Deck & Batygin, 2015; Batygin, 2015; Pichierri et al., 2018; Batygin & Petit, 2023). Nevertheless, most studies focus on one aspect of these issues at a time. Here, we provide a unified framework for studying the dynamics of both resonance capture and subsequent stability. This analysis is applicable to planet pair systems with arbitrary planet mass ratios and at different first-order mean motion resonances. An illustration of this framework can be found in Figure 6. The results are presented in a $\tau_{\rm m}/\tau_{e}$$\tau_{\rm e}$ parameter space, which links all dynamical regimes to the two key parameters that govern planet-disk interactions: angular momentum and eccentricity damping. We describe below how the variations in $\tau_{\rm m}$ and $\tau_{\rm e}$ influence the outcome.
@@
If convergent migration occurs too rapidly, the planets directly bypass the resonance without reaching the equilibrium state. As migration slows down, the effect of eccentricity damping becomes significant. Strong e-damping has two effects. First, it causes a reduction in planet’s eccentricity, weakening the resonant interaction, since the resonant term is tied to planet’s eccentricity ((3)). Furthermore, the semi-major axis also decays due to orbital circularization. If the magnitude of such orbital decay term surpasses the resonant term, the planets will no longer remain in resonance. In this regard, planets cannot be captured when $\tau_{e}$ and/or $\tau_{\rm m}$ are short. No trap occurs in the upper and left parts of Figure 6.
@@
When planets are trapped in resonance, the long-term stability of the system is related to the degree of eccentricity excitation. The damping effect ($\tau_{e}$) reduces the planet’s eccentricity, making the system stable. On the other hand, convergent migration ($\tau_{\rm m}$) brings planets closer together, in turn exciting their eccentricities. Consequently, the relative eccentricity damping $\tau_{\rm m}/\tau_{e}$ is the key parameter in determining the stability. As can be seen from the top to the bottom of Figure 6, the system tends to be more unstable as $\tau_{\rm m}/\tau_{e}$ decreases.